OpenStudy (anonymous):

How do I solve the following equation? 3-2t=4t

OpenStudy (anonymous):

for t? if its for t then its supposed to be -2t-4t= -3 -6t= -3 t= 1/2

OpenStudy (anonymous):

ohhh yah perfect thanks! @jellybean11122

OpenStudy (anonymous):

yea np!!! oh and sorry but i really want to save up on the medals...... do u think u can give me one? its ok if u don't want to :D